Home‎ > ‎Reports‎ > ‎crew110b‎ > ‎

day03


Crew #110B Log Book for January 9, 2012

Photos of the day





Commander's Check-In Report

Date and time: 1/9/2012; 17:00

Written by: April Andreas, Crew 110B, Crew Commander

Crew Physical Status: Good, one crew member had a very slight burn handling a hot shelf from the oven, but it was minor and was treated.

Time Departed/Returned from EVA: 14:00 - 16:30


Brief Narrative of Field Mission Results: We documented the EVAs taken the previous days, with GPS waypoints and pictures, and listed specific tasks for students to do at each waypoint.

EVA Data/Interpretations: None

Engineering/Hab Maintenance: The front bag on ATV 2 has a hole in it. (We discovered this when several critical items went missing and were later found in a variety of places)

There is an AC adapter that's the right voltage for the balance, but the polarity is reversed on it. The balance needs 12V DC, 500 mA, coaxial connector, the inner conductor is negative and the outer conductor is positive.

We need to dispose of the fluorescent light bulbs properly, since they are hazardous waste (we have green-tipped and silver-tipped, the latter are HHW)

We used all the spare fluorescent bulb replacing the burned-out ones in the lab.

Older issues:

ATV 2 still has an oddly-wobbling tire and has a strange smell
The ATV still has a broken tail light
The front plastic on helmet 5 is still broken
Backpack #6 has a loose connection)

Report Transmission Schedule: As available.

Plans for Tomorrow: Tomorrow we will go through the Geology EVA and then begin documentation for that EVA.

Inventory: n/a

Support Requested: See engineering / Hab maintenance

Miscellaneous: n/a

Commander's Report

Date and time: 9 January 2012, 19:31
Written by: April Andreas, Crew 110B, Crew Commander

This was an absolutely fantastic day. We enjoyed a breakfast of leftover pancakes and some microwave bacon and then enjoyed a microbiology lecture from Liz about Gram positive and Gram negative and environmental sampling. We then went downstairs to the lab and learned to create agar media, Maconkey's and Mannitol Salt.

This was an adventure in an of itself, as we had to use an IKEA food scale to get the measurements we needed, and due to the small quantities of material we ended up having to do a negative measure - putting the entire container of media on the scale and weighing what we took out. Next, Larry got to fix a thermocouple so we could take a decent temperature reading on the incubator. After preparing the media, we then poured it into the Petri dishes. Next we all chose specific locations in the hab to swab, completed our swabs, and finally streaked the agar plates.

After a lunch of more leftovers, we split into two groups. Liz stayed in the hab to continue to prepare the plates, Vernon and Larry went to practice working the telescope, and Elaine, Nancy, and I went on an out-of-sim EVA to document the first "In Sim Orientation EVA" that our students will complete when they come out to MDRS. It was necessary to be out of sim because we needed to carefully document every step. Also, the light was fading and we needed to move quickly.

For dinner, Vernon and I prepared a meal of spaghetti and (burned) Alfredo sauce. We also rehydrated green beans and broccoli. Nobody starved, though, particularly since Larry and Liz prepared homemade bread and cherry tarts! Yum!

Tonight, the crew will be going out to the telescope. Tomorrow we will do the Geology EVA and work on the documentation for that.

EVA Report

Date and time: 9 January, 2012

Written by: Elaine Fagner, Crew #110b; Crew Position: Geologist/HSO

Time: 13:48 - 16:15

Crew Members: Elaine Fagner, April Andreas, and Nancy Ray-Mitchell

Site Location: 4.5 miles north of HAB at multiple stops.

Transit Mode: ATV (out of SIM)

Objectives: Development of orientation geology stops with text, assignments, and explanations/pictures for students arriving at the HAB. The 2nd EVA group yesterday did not have a geologist and found it difficult to interpret what they were seeing. They requested that we develop a specific orientation for geology for our students that promoted hands-on experience in SIM to expose them to basic geologic principles, formations, structures, and resources of the area. We designed a 5 stop program that allows students to understand geologic formations, differences between coarse and fine sedimentary rocks, exposure to specific types of sedimentary rocks present in the area, evidence of fluvial deposition, evidence of eolian enviornments, collection of gypsum, petrified wood, and a salt flat.

Results: Team fulfilled mission and developed a full scope orientation that should last no longer than 2 hours from start to finish with students in SIM. It took us longer to develop since we were creating the text, instructions, and experience.

Lessons Learned: This approach was the result of a lesson learned yesterday. We had no issues today and found the approach useful for development of future student curriculum.


Science Report

Date and Time: 9 January, 18:00
Written by: Elizabeth Mitchell, Crew 110B

All members participated this morning in the microbiological assaying of the Hab.

1) All members chose 4 sites within the Hab to swab. Examples include inside the SIM helments and bedroom doors.

2) All members particpated in the making of agar media. We had problems with the scale. We did not have the plug for the scientific scale. We had to use an IKEA food scale. The scale did not measure below grams (did not measure tenths, etc.). For our purposes, this was sufficient. However, the scale was not sensitive enough to pick up small gram amounts of agar. Instead, we had to "negatively weigh" the agar. The entire agar bottle was placed on the scale. We took out the number of grams needed. We also did not have a specific enough way to measure water. We added the necessary mls of water by using 1 g of water = 1 ml. As such, water was weighed.

3) Agar was sterilized in the oven at 150 C for 20 minutes.

4) All members particpated in the pouring of agar plates. All members particpated in the streaking of agar plates.

5) Plates are incubating in the oven at 37 C. We will see what grows tomorrow or the next day, depending on growth activity.


Engineering Report

Date and time: 9 January 2012, 20:00
Written by Nancy Ray-Mitchell, Crew 110B, Chief Engineer

Generator/Electricity:
Did you have any electricity outage? (Yes/No): NO

Fuel Status:
Diesel (Full/0.75/0.5/0.25/Empty): .75
Propane Tank (Full/0.75/0.5/0.25/Empty): .56%
Gasoline Tank (Full/0.75/0.5/0.25/Empty): .25%

Oil:
ATV Oil Quantity (# of Quarts in storage): 2
Car Oil Quantity (# of Quarts in storage): 1

Water Status:
Outside Potable Water Tank Level (inches from bottom): 22
Trailer Potable Water Tank Level (inches from bottom): 2
Hab Potable Water Tank Level (inches from bottom): 27
Potable Water Meter Reading: 489522
Into GreenHab Meter Reading: N/A
Notes/Comments:

GreenHab:
Condition of plants in Tank 1: N/A
Condition of plants in Tank 2: N/A
Condition of plants in Tank 3: N/A

Observatory:
Telescope Used? (Yes/No): YES, was tested to use later tonight.
Observatory Notes/Comments: N/A

Transportation:
Hab Car (NOT RENTAL)
Used (yes/no): YES, by Don
Oil Checked (yes/no): NO
Tire Status: Good
Odometer at end of day: 125,608.5
Notes/Comments on Hab Car: Don took to get inspected today; he is still waiting for the registration.

ATV 1 (Opportunity):
Used: YES
Oil Checked: YES
Fuel Consumed: .33 gal
Tire Status: Good - 4psi
Today's Run Time: 02:00
Season's Run Time: 47:50
Comments:

ATV 2 (Spirit):
Used: YES
Oil Checked: YES
Fuel Consumed: .33 gal
Tire Status: Good - 4psi
Today's Run Time: 02:00
Season's Run Time: 39:53
Comments: It smells like it is burning oil; however, the oil level was fine.

ATV 3 (Viking I):
Used: YES
Oil Checked: YES
Fuel Consumed: .33 gal
Tire Status: Good
Today's Run Time: 02:00
Season's Run Time: 44:35
Comments:

Heating and Ventilation:
Thermostat setting upstairs (Fahrenheit): 76
Actual temperature upstairs (Fahrenheit): 74.8 F

Thermostat setting downstairs (Fahrenheit): 
Downstairs thermostat not working
Actual temperature downstairs (Fahrenheit): 63

Computers/Networking Infrastructure:
Notes/comments on internet/computers?: Internet is sporadic

Journalist Report




Other